Please select To the mobile version | Continue to access the desktop computer version

Electronic Engineer Discuss

12Next
Return to list New
View: 179|Reply: 12

Hantek 6254BD faulty readings and calibration does not end

[Copy link]

1

Threads

10

Posts

10

Credits

新手上路

Rank: 1

Credits
10
Post time 2019-11-17 08:34:09 | Show all posts |Read mode
Hello,

I am using a Hantek 6254BD on a PC running Windows 7.

Yesterday my oscilloscope was working properly. But now it has a problem.


When testing on the 1kHz 2Vpp signal generator, I get a wrong signal. Vpp is meassured much lower (20mv), the Voffset is not right and the meassured signal is distorted.

I tested the signal generator with another oscilloscope and the signal generator is working fine.


I tried different probes and different channels. Same problem.

I tried to calibrate, but the software is stuck in calibration process and the calibration will not end.


Can you please help me.

Kind regards,
Mika



Reply

Use magic Report

28

Threads

1517

Posts

1678

Credits

管理员

Rank: 9Rank: 9Rank: 9

Credits
1678
Post time 2019-11-18 15:57:51 | Show all posts
Please upload a video or picture to study.

1

Threads

10

Posts

10

Credits

新手上路

Rank: 1

Credits
10
 Author| Post time 2019-11-19 03:36:02 | Show all posts
Hello,

I have attached four images to this post to clarify the problem.
Please consider the image captions.

When trying to calibrate the oscilloscope, the calibration process does not make any progress. It stays like it is shown in the image.

Readings have a big voltage offset. I cannot meassure the square wave generator properly, because of voltage offset and the scaling is also wrong.

Thank you in advance for your effort.

Kind regards,
Mika



This post contains more resources

You have to Login for download or view attachment(s). No Account? Register

x

1

Threads

10

Posts

10

Credits

新手上路

Rank: 1

Credits
10
 Author| Post time 2019-11-19 03:50:30 | Show all posts
Some additional information:
Software Version: 20160628
Driver Version: 20160531
FPGA Version: D001
PCB Version: P6.11

28

Threads

1517

Posts

1678

Credits

管理员

Rank: 9Rank: 9Rank: 9

Credits
1678
Post time 2019-11-19 11:16:51 | Show all posts
Please test the voltage at the marked points.



This post contains more resources

You have to Login for download or view attachment(s). No Account? Register

x

1

Threads

10

Posts

10

Credits

新手上路

Rank: 1

Credits
10
 Author| Post time 2019-11-19 17:03:04 | Show all posts
Hello Amy,

I meassured the testing points with following results:
TP10 = -4,387 V
TP4 = 4,622 V

The more interesting thing I found out was, that the oscilloscope is working when it is not in the case.
A bare board without the case works fine.
When I put the board back in the case, it stops working.

In the end I shortened the spring on the back of the board a bit, because I believed it was putting too much pressure on the board.
Now the oscilloscope is working.

Thank you for your help.

Kind regards,
Mika

1

Threads

10

Posts

10

Credits

新手上路

Rank: 1

Credits
10
 Author| Post time 2019-11-19 18:06:38 | Show all posts
Hello Amy,

Because the oscilloscope was not working because of mechanical pressure on the board:
Do you think it is a good idea to repair some solder joints?

Kind regards,
Mika

1

Threads

10

Posts

10

Credits

新手上路

Rank: 1

Credits
10
 Author| Post time 2019-11-21 02:03:54 | Show all posts
Hello,

today the problem came back.
I think it is a broken solder joint, because yesterday the oscilloscope was in a warm environment, now it is about 10°C colder...
I have a feeling that the problem depends on the pressure on the board and on temperature, so I guess it is a broken solder joint...

Can you give me a hint where to look?

Kind regards,
Mika

1

Threads

10

Posts

10

Credits

新手上路

Rank: 1

Credits
10
 Author| Post time 2019-11-21 02:26:25 | Show all posts
Hello,

I warmed the oscilloscope up by bringing it to a warmer room. And there it was working.
Do you have a suggestion how to fix the problem?

Kind regards,
Mika

1

Threads

10

Posts

10

Credits

新手上路

Rank: 1

Credits
10
 Author| Post time 2019-11-21 02:51:02 | Show all posts
When the oscilloscope is not working:
TP10 = +0,335 V
TP4 = +4,632 V
12Next
Return to list New
You have to log in before you can reply Login | Register

Points Rules

QQ|Dark room|Mobile|Archiver|Electronic Engineer Discuss

2019-12-14 05:23 GMT+8 , Processed in 0.482734 second(s), 22 queries .

Powered by Discuz! X3.2

© 2001-2013 Comsenz Inc.

Quick Reply To Top Return to the list